find the length of the altitude to the 26in side of a triangle whose sides are 10,24, and 26 inches in length.
Answer by edjones(8007)

10^2+24^2=26^2 so its a right triangle.
A=.5ab
.5*24*10=120 in^2 area
.
Assuming that the 26" side is the base, x is the altitude.
.5x*26=120
13x=120
x=120/13 in.
.
Ed